# A Survey on Deep Learning Hardware Accelerators for Heterogeneous HPC Platforms

### v2

@@ 5.1. Processing-in-Memory
PIM categories (2) and (3) are mostly implemented on 3D-stacked memory blocks, which rely on the possibility of stacking layers of conventional 2D DRAM or other memory types together with one or more optional layers of logic circuits. These logic layers are often implemented with different process technologies and can include buffer circuitry, test logic, and processing elements. Two main 3D stacked memory standards have been recently proposed: the Hybrid Memory Cube (HMC) and the High Bandwidth Memory (HBM). They both provide highly parallel access to the memory, a sought-after characteristic in the highly parallel architecture of the DNN accelerators. The processing elements of 3D stacked DNN accelerators can be embedded in the logic die or in the memory dies, significantly reducing the latency of accessing data in main memory, and improving the system energy efficiency. However, as detailed in Section A.7 of Appendix A, there are some challenges and limitations to consider when using this technology (Kim2022, ).
@@
Most in-subarray PIMs for DNNs rely on solutions similar to Ambit(Seshadri_2017, ) and RowClone(Seshadri_2013, ) for implementing the computing elements. Ambit is an Accelerator-in-Memory that exploits the analog operation of DRAM technology to perform bit-wise operations completely inside the DRAM. By simultaneously activating three DRAM rows sharing the same set of sense amplifiers, it enables the system to perform bit-wise AND and OR operations. It is also able to perform bit-wise NOT operations by using the inverters available inside slightly modified DRAM sense amplifiers. RowClone is a mechanism that efficiently copies rows inside the same DRAM subarray by exploiting the vast internal DRAM bandwidth without CPU intervention.
@@
DRISA(Li_2017, ) leverages these technologies by implementing bit-wise XORs as basic boolean operations, and by expressing more complex functions as sequences of the basic operation. Additional logic for supporting the calculation (e.g., shifters) is needed, as well as modifications in the memory controllers for driving the execution of operation opcodes. Calculations can be executed with different numbers of bits, but when the number of bits increases the execution time increases exponentially. However, the simultaneous activation of multiple subarrays and banks provides large parallelism and large computational throughput. While DRISA evaluates the implementation of CNNs with binary weights, DrAcc(Deng_2018, ) focuses on CNNs with ternary weights.
@@
The Micron In-Memory Intelligence (IMI) architecture(Finkbeiner_2017, ) is built on simple bit-serial computing elements placed below standard DRAM array’s sense-amplifiers and provides the memory block with the ability for massive SIMD parallelism by supporting vector instructions over an entire bank. The simple circuits can execute basic logical functions (such as XOR and AND) and implement more complex functions as a serial sequence of them. A control unit is attached to each DRAM bank and translates the IMI instruction to be executed into row-cycles that control the SIMD computing elements. Data can be organized either horizontally or vertically, and the data layout within a subarray can be changed with minimum penalty.
@@
SIMDRAM(Hajinazar_2021, ) is a flexible general-purpose processing-using-DRAM framework that enables the efficient implementation of complex operations and provides a flexible mechanism to support the implementation of arbitrary user-defined operations. SIMDRAM makes use of a subarray organization that incorporates additional functionality to perform logic primitives (i.e., MAJ and NOT) and uses a sequence of these basic operations to build more complex functions. The proposed framework generates the required sequence of DRAM commands to perform the desired operation in DRAM, which are executed by a control unit located inside the memory controller which manages the computation of the operation from start to end.
@@
PIM-DRAM(Roy_2021, ) is another example of in-subarray PIM, which, differently from previous proposals, performs multiplication operations at the subarray level, while the accumulation and activation functions are performed in the bank architecture. The multiplication is performed as a sequence of bit-wise in-subarray addition and AND operations which are implemented by dedicating extra ”compute rows” in each sub-array for copying the original data and storing intermediate computational values. The processing happens by enabling multiple wordlines at the same time to leverage the large available internal DRAM bandwidth. As in previous work, PIM-DRAM relies on the RowClone to transfer data among banks, and a reconfigurable adder tree is added to the peripheral area for intra-bank accumulation.
@@
{forest}
SIMDRAM(Hajinazar_2021, ), PIM-DRAM(Roy_2021, )
] ] [Logic die (3D-DRAM) [ Neurocube(Kim2016, ), Tetris(Gao2017, )
NeuralHMC(Min2019, ), VIMA(Cordeiro2021, )
] ] [Bank (3D-DRAM) [ Newton(He2020, )
HBM-PIM(Kwon2021, )
] ] ], [IMC [SRAM [Orlando (desoli17isscc, ),
STM-IMC (desoli2023isscc, ) ] ] [PCM [ Hermes (khaddam2022hermes, ),
NeuralPIM (cao2022tcomp, ), PRIME (chi2016isca, ), NeuRRAM (wan2022nature, )
@@
A first example of a logic die-level PIM implementation is Neurocube (Kim2016, ) that, as shown in Figure 7(a), is embedded into the logic die of an HMC, and consists of a cluster of PEs connected by a 2D mesh Network-on-Chip (NoC). The PE is composed of a row of multiply accumulator (MAC) units, a cache memory, a temporal buffer, and a memory module for storing shared synaptic weights. Each PE is associated with a single memory vault and can operate independently and communicate through the TSVs and the vault controller. A host communicates with the Neurocube through the external links of the HMC to configure the Neurocube for different neural network architectures. Each vault controller in the HMC has an associated programmable neuro sequence generator (PNG), i.e., a programmable state machine that controls the data movements required for neural computation. Neurocube implements an output stationary dataflow, meaning that each MAC from a PE is responsible for the computations of a different output neuron at a time.
@@
Similarly to Neurocube, Tetris (Gao2017, ) uses an HMC memory stack organized into 16 vaults. Each vault is associated with a PE, connected to the vault controller, and composed of a systolic array of 14 $\times$14 PEs and a small SRAM buffer, shared among the PEs. A 2D mesh NoC connects all the PEs. Differently from previous accelerator approaches, the dimension of the buffers in the logic layer is reduced and optimized to take into account the lower cost of accessing the DRAM layers, as well as the area constraints of the 3D package. Each PE has a register file and a MAC locally storing the inputs/weights and performing computations. Tetris implements a row stationary dataflow that maps 1D convolutions onto a single PE and utilizes the PE register file for local data reuse. A 2D convolution is orchestrated on the 2D array interconnect so that the data propagation among PEs remains local. In (Gao2017, ), an optimal scheduling is discussed to maximize on-chip reuse of weights and/or activations, and resource utilization. However, a programming model is not presented.
@@
NeuralHMC (Min2019, ) adopts the same systolic architecture and row-stationary dataflow discussed in Tetris. However, the authors introduce a weight-sharing pipelined MAC design to lower the cost of accessing weight data, by reducing the original 32-bit floating-point weights to a 5 or 8-bit cluster index, saving memory consumption. Moreover, they discuss a series of mechanisms to reduce and optimize packet scheduling and on-chip communication in multi-HMC architectures.
@@
The authors in (Cordeiro2021, ) study the benefits of migrating ML kernels on near-data processing (NDP) architecture capable of large-vector operations. The work derives from previous work of the same authors (Alves2016, ), where they introduced the HIVE architecture, which extends the HMC ISA for performing common vector operations directly inside the HMC, avoiding contention on the interconnections as well as cache pollution. The newly introduced Vector-In-Memory Architecture (VIMA) supports all ARM NEON Integer and floating-point instructions, operating over vectors of 8 KB of data by fetching data over the 32 channels (vaults) of the HMC in parallel. The authors extend and use an NDP intrinsics library that supports the validation of NDP architectures based on large vectors, provides insights, and show the benefits of migrating ML algorithms to vector-based NDP architectures. Their simulated results show a significant speed-up and energy reduction with respect to an x86 baseline.
@@
Several accelerators adopting the bank-level PIM approach can be found in the literature. Newton (He2020, ) proposes a fixed data flow accelerator that computes matrix-vector multiplication effectively which is implemented at the bank level of an HBM. It employs only MAC units, buffers, and a DRAM-like command interface for the host CPU to issue commands to the PIM compute, avoiding the overhead and granularity issues of offloading-based accelerators, e.g., the delay in the launch of the kernel and the switching between the PIM/non-PIM operational modes. To reduce the output vector write traffic with minimal output buffering, Newton employs an unusually wide interleaved layout (DRAM row-wide). In Newton, input/output vectors have high reuse while the matrix has no reuse.
@@
HBM-PIM (Kwon2021, ) implements a function-in-memory DRAM (FIMDRAM) that integrates a 16-wide SIMD engine within the memory banks exploits bank-level parallelism to provide 4 $\times$ higher processing bandwidth than an off-chip memory solution (Figure 7(b)). In their design, half of the cell array in each bank of the HBM was removed and replaced by a programmable computing unit (PCU), placed adjacent to the cell array to utilize bank-level parallelism. Each PCU is shared among two banks, and there are 8 PCUs per pseudo-channel. The PCU is divided into a register group, an execution unit, a decoding unit for parsing instructions needed to perform operations, and interface units to control data flow. The register group consists of a command-register file for instruction memory (CRF), a general-purpose register file for weight and accumulation (GRF), and a scalar register file to store constants for MAC operations (SRF). The PIM controller is integrated to support the programmability of the PCU and, similarly to Newton, the seamless integration with the host by determining the switching between the PIM/non-PIM operational modes. If the PIM mode is asserted, the PCUs execute the instructions pre-stored in the CRF, incrementing the program counter every time a DRAM’s read command is issued. 3D-stacked PIM has also been proposed for accelerating applications loosely related to DNNs. We present a brief overview of these accelerators in Section A.7 of the Appendix A.
@@ 5.2. In-Memory Computing
In-memory computing (IMC) has been proposed to break both the memory and the compute wall in data-driven AI workloads, using either SRAM or emerging memory technologies (such as PCM and RRAM described in Section A.8 of Appendix A) offering different trade-offs when used as an integrated computing device at the system level.
Full-digital IMC designs offer a fast path for the integration of the next generation of neural processing systems like NPUs. An example of IMC architecture has recently been proposed by STMicroelectronics in (desoli2023isscc, ), where a scalable and design time parametric NPU for edge AI relying on digital SRAM IMC has been manufactured in 18 nm FDSOI technology achieving an end-to-end system-level energy efficiency of 77 TOPS/W and an area efficiency of 13.6 TOPS/mm2. This IMC-NPU is the evolution of the Orlando system-on-chip proposed in (desoli17isscc, ) and is specialized in accelerating the inference workloads. Another digital IMC design is NeuroCIM (kim2022vlsi, ), an energy-efficient processor with four key features achieving 310.4 TOPS/W using dedicated hardware design to lower the activity of the memory early terminating the operations when needed.
@@
A first example of a non-volatile inference-based machine on RRAM technology is the ISAAC (shafiee2016isca, ) tile-based architecture that proposes a pipeline design for CNN processing which combines the data encoding and the processing steps within in situ multiply and accumulate units (IMA). The design is pipelined fetching data from an external eDRAM chip to the computing tile. The data format in ISAAC is fixed at 16-bit. During computation, at each clock cycle, 1-bit is given as input to the IMA, and then the result from the IMA is converted to the digital format, thus requiring 16 clock cycles to process the input. Such a design allows implementing the computation on different tiles in a fully pipelined approach to increase computing performance. Tiled computation is widely used from RRAM to improve the throughput. The PipeLayer (song2017hpca, ) architecture introduces intra-layer parallelism and an inter-layer pipeline for tiled architecture, using duplicates of processing units featuring the same weights to process multiple data in parallel. Designs like PRIME (chi2016isca, ) take part of the RRAM memory arrays to serve as acceleration instead of adding an extra processing unit for computation. The work in (cao2022tcomp, ) outlines that existing PIM RRAM accelerators suffer from frequent and energy-intensive analog-to-digital (A/D) conversions, severely limiting their performance. To this extent, they presented a new architecture to efficiently accelerate DL tasks by minimizing the required A/D conversions with analog accumulation and neural-approximated peripheral circuits. Using a new dataflow, they remarkably reduced the required A/D conversions for matrix-vector multiplications by extending shift-and-add operations to the analog domain before the final quantization. A summary of the technological features in major RRAM accelerators can be found in (smagulova2023pieee, ).
@@
The first PCM-based silicon demonstrator for DNN inference is Hermes, which appeared in 2021 (khaddam2022hermes, ). The IMC accelerator consists of a 256x256 PCM cross-bar and optimized ADC circuitry to reduce the read-out latency and energy penalty. The SoC is implemented in 14nm technology, showing 10.5 TOPS/W energy efficiency and performance density of 1.59 TOPS/mm2. The same 256x256 PCM cross-bar has been integrated into a scaled-up mixed-signal architecture that targets the inference of long short-term memory (LSTM) and ResNet-based neural networks (gallo202264, ). The chip, implemented in the same 14nm technology, consists of 64 analog cores interconnected via an on-chip communication network and complemented with digital logic to execute activation functions, normalization, and other kernels than Matrix-Vector Multiplications (MVMs). The accelerator achieves a peak throughput of 63.1 TOPS with an energy efficiency of 9.76 TOPS/W for 8-bit input/8-bit output MVM operations.
@@ 5.3. Neuromorphic Accelerators
A first example of a digital architecture for SNN and neuroscience simulation acceleration is the SpiNNaker chip (painkras2013jsscc, ). It follows a distributed von Neumann approach using a globally asynchronous locally synchronous (GALS) design for efficient handling of asynchronous spike data and is based on a 130 nm technology. SpiNNaker has been optimized for large-scale SNN experiments while keeping a high degree of flexibility. The evolution of the architecture using 22 nm technology embedding 4 ARM Cortex M4F cores out of the 152 per chip is planned for the final SpiNNaker 2 system (liu2018frontiers, ). The objective is to simulate two orders of magnitude more neurons per chip compared to (painkras2013jsscc, ). However, it has been demonstrated that GPU-based accelerators compare favorably to a SpiNNaker system when it comes to large SNN and cortical-scale simulations (knight2018frontiers, ).
@@
Full-custom digital hardware leads to higher-density and more energy-efficient neuron and synapse integration for spiking neural networks (SNN) compared to the two formerly described accelerators (frenkel2021arxiv, ). All the accelerators reported in this paper benefit from moving the memory (generally SRAM elements) closer to computation. The 45 nm design in (seo2011cicc, ) is a small-scale architecture for SNN acceleration embedding 256 Leaky-Integration-Fire (LIF) neurons and up to 64k synapses based on the Stochastic Synaptic Time Dependant Plasticity (S-STDP) concept. It achieves reasonably high neuron and synapse densities, despite the use of a custom SRAM and given its energy-efficiency figures is an ideal choice, especially for edge computing scenarios. At the same integration scale, the ODIN chip embeds 256 neurons and 64k Spike Driven Synaptic Plasticity (SDSP)-based 4-bit synapses in a 28 nm CMOS process (frenkel2018tbiocas, ). A first attempt to scale up the NPU for SNN applications is represented by the 65 nm MorphIC chip, which is based on the ODIN core integrated into a quadcore design (frenkel2019tbiocas, ).
@@
Concerning large-scale neuromorphic platforms required for cognitive computing applications, there are currently two designs offered: the 28 nm IBM TrueNorth (akopyan2015tcad, ) and the 14 nm Intel Loihi (davies2018micro, ) neuromorphic chips. TrueNorth is a GALS design embedding as high as 1M neurons and 256M binary non-plastic synapses per chip, where neurons rely on a custom model that allows modifying their behaviors by combining up to three neurons (cassidy2013ijcnn, ). Loihi is a fully asynchronous design embedding up to 180k neurons and 114k (9-bit) to 1M (binary) synapses per chip. Neurons rely on a LIF model with a configurable number of compartments to which several functionalities such as axonal and refractory delays, spike latency, and threshold adaptation have been added. The spike-based plasticity rule used for synapses is programmable. Loihi will evolve then in the new Loihi 2 neuromorphic chip and TrueNorth into the NorthPole platform (northpole, ).
@@
In digital designs for neuromorphic chips, versatility can be obtained with a joint optimization comprising power and area efficiencies. This flexibility in optimizing between versatility and efficiency in digital designs is highlighted with platforms going from versatility-driven (e.g., SpiNNaker) to efficiency-driven (e.g., ODIN and MorphIC), through platforms aiming at a well-balanced trade-off on both sides (e.g., Loihi). Table 8 summarizes the main characteristics of the neuromorphic chips described so far with particular insight on the Energy per spike operation (SOP) that is seen as a primary benchmarking factor for these architectures.

### v3

@@ 5.1. Processing-in-Memory
3D-stacked memory blocks rely on the possibility of stacking layers of conventional 2D DRAM or other memory types together with one or more optional layers of logic circuits. These logic layers are often implemented with different process technologies and can include buffer circuitry, test logic, and PEs. Two main 3D stacked memory standards have been recently proposed: the Hybrid Memory Cube (HMC) and the High Bandwidth Memory (HBM). They both provide highly parallel access to the memory, a sought-after characteristic in the highly parallel architecture of the DNN accelerators. The PEs of 3D stacked DNN accelerators can be embedded in the logic die or in the memory dies, significantly reducing the latency of accessing data in main memory, and improving the system energy efficiency. However, as detailed in Section A.7 of Appendix A, there are some challenges and limitations to consider when using this technology (Kim2022, ).
@@
Most in-subarray PIMs for DNNs rely on solutions similar to Ambit (Seshadri_2017, ) and RowClone (Seshadri_2013, ) for implementing the computing elements. Ambit exploits the analog operation of DRAM technology to perform bit-wise AND, OR and NOT operations completely inside the DRAM. RowClone is a mechanism that efficiently copies rows inside the same DRAM subarray by exploiting the vast internal DRAM bandwidth without CPU intervention.
@@
DRISA(Li_2017, ) leverages these technologies by implementing bit-wise XORs, and by expressing more complex functions as sequences of such a basic operation. Additional logic (e.g., shifters) and modifications in the memory controllers are needed for driving the execution of operation opcodes. Higher bit-widths are supported, but with the execution time increasing exponentially. However, multiple subarrays and banks provide large parallelism and large computational throughput. While DRISA evaluates the implementation of CNNs with binary weights, DrAcc (Deng_2018, ) focuses on CNNs with ternary weights.
@@
The Micron In-Memory Intelligence (IMI) architecture (Finkbeiner_2017, ) is built on simple bit-serial computing elements placed below standard DRAM array’s sense-amplifiers and provides the memory block with the ability for massive SIMD parallelism by supporting vector instructions over an entire bank. Complex operations are implemented as serial sequences of basic logic functions, such as XOR and AND. A control unit is attached to each DRAM bank and translates the IMI instruction to be executed into row-cycles that control the SIMD computing elements.
@@
SIMDRAM (Hajinazar_2021, ) is a flexible general-purpose processing-using-DRAM framework that provides a flexible mechanism to support the implementation of arbitrary user-defined operations as sequences of basic functions including MAJ and NOT. The sequence of DRAM commands generated by the framework are executed by a control unit located inside the memory controller, which manages the computation from start to end.
@@
PIM-DRAM(Roy_2021, ) is an in-subarray PIM, which performs multiplications as sequences of bit-wise in-subarray addition and AND operations while the accumulation and activation functions are performed in the bank architecture. The processing happens by enabling multiple wordlines at the same time to leverage the large available internal DRAM bandwidth.
@@
A first example of a logic die-level PIM implementation is Neurocube (Kim2016, ) that, as shown in Figure 7(a), is embedded into the logic die of an HMC, and consists of a cluster of PEs connected by a 2D mesh Network-on-Chip (NoC). The PE is composed of a row of multiply accumulator (MAC) units, a cache memory, a temporal buffer, and a memory module for storing shared synaptic weights. Each PE is associated with a single memory vault and can operate independently and communicate through the TSVs and the vault controller. A host communicates with the Neurocube through the external links of the HMC to configure the Neurocube for different neural network architectures. Each vault controller in the HMC has an associated programmable neuro sequence generator (PNG), i.e., a programmable state machine that controls the data movements required for neural computation. Neurocube implements an output stationary dataflow.
@@
Tetris (Gao2017, ) uses an HMC memory stack organized into 16 vaults. Each vault is associated with a PE, connected to the vault controller, and composed of a systolic array of 14 $\times$14 PEs and a small SRAM buffer, shared among the PEs. A 2D mesh NoC connects all the PEs. The dimension of the buffers in the logic layer is reduced and optimized to take into account the lower cost of accessing the DRAM layers, as well as the area constraints of the 3D package. Each PE has a register file and a MAC locally storing the inputs/weights and performing computations. Tetris implements a row stationary dataflow that maps 1D convolutions onto a single PE.A 2D convolution is orchestrated on the 2D array interconnect so that the data propagation among PEs remains local. In (Gao2017, ), an optimal scheduling is discussed to maximize on-chip reuse of weights and/or activations, and resource utilization. However, a programming model is not presented.
@@
NeuralHMC (Min2019, ) adopts a weight-sharing pipelined MAC to lower the cost of accessing weight data, by reducing the original 32-bit floating-point weights to a 5 or 8-bit cluster index, saving memory consumption. Moreover, it allows reducing and optimizing packet scheduling and on-chip communication in multi-HMC architectures.
@@
The HIVE architecture (Alves2016, ) extends the HMC ISA for performing common vector operations directly inside the HMC. By migrating ML kernels on near-data processing (NDP) architectures capable of large-vector operations, the Vector-In-Memory Architecture (VIMA) proposed in (Cordeiro2021, ) supports all ARM NEON Integer and floating-point instructions, operating over vectors of 8 KB of data by fetching data over the 32 channels (vaults) of the HMC in parallel. In this way, it leads to a significant speed-up and energy reduction with respect to an x86 baseline.
@@
Several accelerators adopting the bank-level PIM approach can be found in the literature. The Newton fixed data flow accelerator proposed in (He2020, ) employs only MAC units, buffers, and a DRAM-like command interface with the host CPU, avoiding the overhead and granularity issues of launching the kernel and switching between the PIM/non-PIM operational modes. The output vector write traffic is reduced by means of an unusually wide interleaved layout (DRAM row-wide). Moreover, input/output vectors have high reuse while the matrix has no reuse.
@@
HBM-PIM (Kwon2021, ) implements a function-in-memory DRAM (FIMDRAM) that integrates a 16-wide SIMD engine within the memory banks exploits bank-level parallelism to provide 4 $\times$ higher processing bandwidth than an off-chip memory solution (Figure 7(b)). Each computing unit (PCU) is shared among two banks, and there are 8 PCUs per pseudo-channel. The PCU is divided into a register group, an execution unit, a decoding unit for parsing instructions needed to perform operations, and interface units to control data flow. The register group consists of a command-register file for instruction memory (CRF), a general-purpose register file for weight and accumulation (GRF), and a scalar register file to store constants for MAC operations (SRF). The PIM controller is integrated to support the programmability of the PCU and the seamless integration with the host by determining the switching between the PIM/non-PIM operational modes. If the PIM mode is asserted, the PCUs execute the instructions pre-stored in the CRF, incrementing the program counter every time a DRAM’s read command is issued. 3D-stacked PIM has also been proposed for accelerating applications loosely related to DNNs. We present a brief overview of these accelerators in Section A.7 of the Appendix A.
@@ 5.2. In-Memory Computing
In-memory computing (IMC) has been proposed to break both the memory and the compute wall in data-driven AI workloads, using either SRAM or emerging memory technologies (such as PCM and RRAM described in Section A.8 of Appendix A) integrated in a dedicated accelerator (Table 7).
@@
Full-digital IMC designs offer a fast path for the integration of the next generation of neural processing systems like NPUs. Recently, STMicroelectronics proposed a scalable and design time parametric IMC-NPU relying on digital SRAM IMC for edge AI (desoli2023isscc, ). This architecture is the evolution of the Orlando SoC (desoli17isscc, ) and is specialized in accelerating the inference workloads. When manufactured in 18 nm FDSOI technology, this IMC-NPU achieves an energy efficiency of 77 TOPS/W and an area efficiency of 13.6 TOPS/mm2. With its four key features and dedicated hardware able to lower the activity of the memory early terminating the operations when needed, NeuroCIM (kim2022vlsi, ) achieves 310.4 TOPS/W. The ISAAC non-volatile inference-based machine on RRAM technology (shafiee2016isca, ) is a tile-based architecture for CNN processing which combines the data encoding and the processing steps within in situ MAC units (IMA). The design is pipelined fetching data from an external eDRAM chip to the computing tile. The data format in ISAAC is fixed at 16-bit. During computation, at each clock cycle, 1-bit is given as input to the IMA, whose result is converted to the digital format, thus requiring 16 clock cycles to process the input. Such a design allows implementing the computation on different tiles in a fully pipelined approach to increase computing performance and throughput. The PipeLayer (song2017hpca, ) architecture introduces intra-layer parallelism and an inter-layer pipeline for tiled architecture, using duplicates of processing units featuring the same weights to process multiple data in parallel. Designs like PRIME (chi2016isca, ) take part of the RRAM memory arrays to serve as acceleration instead of adding an extra processing unit for computation. As outlined in (cao2022tcomp, ), existing PIM RRAM accelerators suffer from frequent and energy-intensive analog-to-digital (A/D) conversions, severely limiting their performance. To efficiently accelerate DL tasks by minimizing the required A/D conversions, a new architecture was presented with analog accumulation and neural-approximated peripheral circuits. The new dataflow introduced in (cao2022tcomp, ) remarkably reduces the required A/D conversions for matrix-vector multiplications by extending shift-and-add operations to the analog domain before the final quantization. A summary of the technological features in major RRAM accelerators can be found in (smagulova2023pieee, ).
@@
The first PCM-based silicon demonstrator for DNN inference is Hermes (khaddam2022hermes, ) which consists of a 256x256 PCM cross-bar and optimized ADC circuitry to reduce the read-out latency and energy penalty. The SoC is implemented in 14nm technology, showing 10.5 TOPS/W energy efficiency and performance density of 1.59 TOPS/mm2. The same 256x256 PCM cross-bar has been integrated into a scaled-up mixed-signal architecture that targets the inference of long short-term memory (LSTM) and ResNet-based neural networks (gallo202264, ). The chip, implemented in the same 14nm technology, consists of 64 analog cores interconnected via an on-chip communication network and complemented with digital logic to execute activation functions, normalization, and other kernels than Matrix-Vector Multiplications (MVMs). The accelerator achieves a peak throughput of 63.1 TOPS with an energy efficiency of 9.76 TOPS/W for 8-bit input/8-bit output MVM operations.
@@ 5.3. Neuromorphic Accelerators
SpiNNaker chip (painkras2013jsscc, ) is a digital architecture designed on a 130 nm technology for SNN and neuroscience simulation acceleration. It is based on a distributed von Neumann approach using a globally asynchronous locally synchronous (GALS) design for efficient handling of asynchronous spike data. The SpiNNaker 2 system (liu2018frontiers, ) uses a 22 nm technology and embeds 4 ARM Cortex M4F cores out of the planned 152 per chip. The objective is to simulate two orders of magnitude more neurons per chip compared to (painkras2013jsscc, ). However, it has been demonstrated that GPU-based accelerators compare favorably to a SpiNNaker system when it comes to large SNN and cortical-scale simulations (knight2018frontiers, ).
@@
In comparison with the above-described accelerators, full-custom digital hardware leads to higher-density and more energy-efficient neuron and synapse integration for SNN (frenkel2021arxiv, ). The 45 nm design in (seo2011cicc, ) is a small-scale architecture embedding 256 Leaky-Integration-Fire (LIF) neurons and up to 64k synapses based on the Stochastic Synaptic Time Dependant Plasticity (S-STDP) concept. Due to its reasonably high neuron and synapse densities and energy-efficiency, this design is an ideal choice for edge computing scenarios. At the same integration scale, the ODIN chip embeds 256 neurons and 64k Spike Driven Synaptic Plasticity (SDSP)-based 4-bit synapses in a 28 nm CMOS process (frenkel2018tbiocas, ). A first attempt to scale up the NPU for SNN is represented by the 65 nm MorphIC chip, which is based on the ODIN core integrated into a quadcore design (frenkel2019tbiocas, ).
@@
Two large-scale neuromorphic platforms required for cognitive computing applications, are currently offered: the 28 nm IBM TrueNorth (akopyan2015tcad, ) and the 14 nm Intel Loihi (davies2018micro, ). TrueNorth is a GALS design embedding as high as 1M neurons and 256M binary non-plastic synapses per chip, where neurons rely on a custom model that allows modifying their behaviors by combining up to three neurons (cassidy2013ijcnn, ). Loihi is a fully asynchronous design embedding up to 180k neurons and 114k (9-bit) to 1M (binary) synapses per chip. Neurons rely on a LIF model with a configurable number of compartments to which several functionalities such as axonal and refractory delays, spike latency, and threshold adaptation have been added. The spike-based plasticity rule used for synapses is programmable. Loihi will evolve then in the new Loihi 2 neuromorphic chip and TrueNorth into the NorthPole platform (northpole, ).
@@
Digital designs for neuromorphic chips can obtain versatility with a joint optimization of power and area efficiencies. This flexibility is highlighted with platforms going from versatility-driven (e.g., SpiNNaker) to efficiency-driven (e.g., ODIN and MorphIC), through platforms aiming at a well-balanced trade-off on both sides (e.g., Loihi). Table 8 summarizes the main characteristics of the neuromorphic chips described so far with particular insight on the Energy per spike operation (SOP).